Can you help me with this problem
It is 2(4-2r)=-2(r+5)

Answer by doukungfoo(195)

Solving for r
2(4-2r)=-2(r+5)
start on the left side of the equal sign and distribute 2 across (4-2r)
This means 2*4 and 2 * -2r
8 - 4r = -2(r+5)
now on the right side of the equal sign distribute -2 across (r+5)
8 - 4r = -2r - 10
move 8 to the right side of the equal sign by subtracting it from both sides
8 - 8 - 4r = -2r - 10 - 8
8's on the left side cancel out leaving
-4r = -2r - 10 - 8
-4r = -2r - 18
next move 2r to the left side by adding it to both sides
-4r + 2r = -2r + 2r - 18
2r's on the right side cancel out leaving
-4r + 2r = -18
-2r = -18
slove for r by dividing both sides by -2
-2r/-2 = -18/-2
r = 9
Check answer by trying it in the orginal equation
2(4-2r)=-2(r+5)
2(4-2(9))=-2(9+5)
2(4-18)=-2(14)
2(-14)=-2(14)
-28 = -28